OpenStudy (anonymous):

how do i simplify (2y^2x^3h)^2

OpenStudy (anonymous):

(2y^2x^3h)^2 is the same as: (2y^2x^3h)*(2y^2x^3h) then you times everything on the left by everything on the right like so: 2y*2y = 4y, 2y*2x = 4yx, 2y*3h = 6yh so far, we now have 4y+4yx+6yh then we go do the same for the rest: 2x*2y = 4xy, 2x*2x = 4x, 2x*3h = 6xh so now we combine like terms, and add up those equations: 4y+4yx+6yh+4xy+4x+6xh now we do the same for the last bit: 3h*2y = 6hy, 3h*2x = 6hx, 3h*3h = 9h now we combine the like terms again, which there is no more: 4y + 4yx + 6th + 4xy + 4x + 6xh + 6hy + 6hx + 9h is now the most simplified form i believe

OpenStudy (anonymous):

oops, not 6th, its 6yh, damn autocorrect